# GravSphere2: A higher-order Jeans method for mass-modeling spherical stellar systems

### v2

@@ Abstract
Aims. Mass-modeling methods are used to infer the gravitational field of stellar systems, from globular clusters to giant elliptical galaxies. While many methods already exist, most require either assumptions about the form of the underlying distribution function or data binning that leads to some loss of information. Furthermore, when only line-of-sight (LOS) data are available, many methods suffer from the well-known mass-anisotropy degeneracy. To overcome these limitations, we develop a new and publicly available mass-modeling method, GravSphere2. This combines individual stellar velocities from LOS and proper motion (PM) measurements to solve the Jeans equations up to fourth order, without any data binning. Using flexible functional forms for the velocity anisotropy profiles at second and fourth order, we show how including additional constraints from a new observable – fourth-order PMs – fully solving along the three dimensions and breaking the mass-anisotropy degeneracy at all orders. We test our method on mock data for dwarf galaxies, showing how GravSphere2 improves on previous methods.
Methods. GravSphere2 introduces four key improvements over previous Jeans mass-modeling methods in the literature: (i) we include fourth-order velocity moment equations in both the LOS and PM directions, for the first time, using these to break model degeneracies; (ii) we use a fully general treatment of both the second and fourth-order velocity anisotropies; (iii) we introduce a ‘bin-free’ approach where we fit individual tracer velocities and positions using flexible and self-consistent probability density functions that include kurtosis; and (iv) we improve the likelihood sampling by using the nested sampler dynesty.
Results. GravSphere2 recovers the mass density, stellar velocity anisotropy, and the logarithmic slope of the mass density profile within its quoted 95% confidence intervals across almost all mocks over a wide radial range ($0.1\lesssim r/R_{1/2}\lesssim 10$, where $R_{1/2}$ is the projected half-light radius). As the number of tracers is lowered – even down to just 10 tracers – it gracefully degrades, with larger uncertainties but no induced bias. We find that GravSphere2 outperforms simple mass estimators, suggesting that it is worth using even when only a few LOS velocities are available. With 1,000 tracers without PMs, GravSphere2 recovers the logarithmic density slope at $R_{1/2}$ with $12\%$ (25%) statistical errors for cuspy (cored) mock data, enabling a distinction between the two. Including PMs, this improves to $8\%$ (12%). With just 100 tracers and no PMs, we are still able to recover slopes with $\sim 30\%$ (20%) errors. GravSphere2 will be a valuable new tool to hunt for massive black holes and invisible dark matter in spherical stellar systems, from globular clusters and dwarf galaxies to giant ellipticals and galaxy clusters.
@@ 1 Introduction
Many stellar systems are approximately spherical, including globular clusters (GCs) and nearby dwarf spheroidal galaxies (dSphs). Both GCs and dSphs contain seemingly invisible mass components that can be inferred from the anomalous motions of their stars and gas (see e.g. 1978PhDT.......195B; 1980ApJ...238..471R; 1987ARA&A..25..425T; 2002MNRAS.333..697L; van2010new; battaglia2013internal, for more recent examples in GCs and dSphs (respectively), see e.g. Evans:2021bsh; 2024MNRAS.529..331D; 2024ApJ...975..268S; 2025A&A...693A.104B and 2019MNRAS.484.1401R; 2020ApJ...904...45H; 2022NatAs...6..659B; 2023A&A...676A..63B; 2024MNRAS.535.1015D; 2025A&A...699A.347A; 2025A&A...700A..77P; 2025arXiv250820711V). For GCs, the invisible mass most likely constitutes stellar remnants or an ‘intermediate-mass black hole’ (IMBH), though some GCs may also contain dark matter. By contrast, dSphs appear to be dominated at all radii by dark matter. With increasingly available spectroscopic and astrometric data from recent surveys, GCs promise an unprecedented natural laboratory to study stellar remnants and hunt for IMBHs (2017IJMPD..2630021M; 2020A&ARv..28....4N; 2020ARA&A..58..257G; 2023arXiv231112118A; 2023MNRAS.522.5740V; 2024Natur.631..285H; 2025A&A...693A.104B), while dSphs promise the same for dark matter (2017arXiv170701348T; 2018MNRAS.481..860R; 2019MNRAS.484.1401R; 2022NatAs...6..659B; 2024MNRAS.535.1015D; 2024ApJ...970....1V; 2025arXiv250820711V; 2025PhRvL.134o1001Z).
The above proliferation in data has led to an associated proliferation in mass-modeling methodologies. All methods seek to solve the steady-state Collisionless Boltzmann Equation (CBE), also known as Vlasov Equation (e.g. Binney2008):
@@
where $\Phi$ is the gravitational potential resulting from all mass components in the system, and satisfying the Poisson equation with mass density $\rho$:
@@
By solving for the gravitational potential and subtracting the expected potential from all visible components we can infer, and study, the presence of invisible masses like black holes and dark matter (e.g. van2010new; 2009ApJ...704.1274W), which is usually done in the steady state approximation, assuming $\frac{\partial f}{\partial t}=0$.
@@
There are number of methods for solving the CBE. Jeans methods solve moments of the CBE (1922MNRAS..82..122J), having the advantages that they are fast [^1] This is particularly timely when assessing increasingly large samples with thousands (in some cases approaching millions, 2025ApJ...983...95H) of tracers. and minimize potential biases that could creep in due to assumptions about the form of the distribution function, $f$ (e.g. 2008MNRAS.390...71C; 2013MNRAS.436.2598W; 2013MNRAS.429.3079M; 2015arXiv150405533C; 2017MNRAS.471.4541R). The main disadvantages, however, are that the data typically need to be binned, while $f$ is not determined. This leads to some loss of information and makes it hard to incorporate stars of uncertain membership, binaries and/or rigorously model selection effects (e.g. 2021MNRAS.501..978R). Furthermore, since $f$ is not determined, a physical $f$ is not guaranteed. This can result in models that formally require a negative $f$ in some regions of phase space (1992ApJ...391..531D; 2013MNRAS.432.3361R) and that should, therefore, be discarded (e.g. 2006ApJ...642..752A).
@@
Distribution function methods assume some form for $f$ and then fit this to the data (e.g. 2015MNRAS.454..576G; 2019MNRAS.482.1525V; 2021MNRAS.501..978R; 2023MNRAS.522.5320D). $f$ can also be implicitly expressed as a superposition of orbits with ‘Schwarzschild methods’ (e.g. 1979ApJ...232..236S; 2008ApJ...682..841C; 2008MNRAS.385..647V; 2013MNRAS.433.3173B) or with ‘made-to-measure’ methods, in which $N$-body simulations are evolved to fit some data constraints (e.g. 1996MNRAS.282..223S; 2007MNRAS.376...71D; 2009MNRAS.395.1079D; 2013MNRAS.430.1928H), while 2025MNRAS.538.1442L propose a novel non-parametric distribution function method. Distribution function methods have the advantages that no data binning is required and $f$ is determined by the data. However, if the functional form for $f$ (or its implicit specification via other methods) does not encompass the true solution, then bias can creep in.
@@
Finally, numerous machine-learning methods have been proposed that learn from training data to either accelerate distribution function fitting or incorporate ‘cosmological priors’ into the mass-modeling pipeline (e.g. 2023MNRAS.519.4384E; 2025arXiv250303812N; 2025arXiv250500763L; 2025arXiv250307717S).
@@
Given the proliferation of mass-modeling methods, we may wonder why yet another method is needed. However, there are two key problems that remain to be addressed, even for spherical stellar systems. Firstly, while the amount and quality of data continue to improve, there will always be interesting systems with very few stellar tracers (e.g. 2021MNRAS.505.5686C; 2024ApJ...961...92S; 2025arXiv250504475P). Typically, simple mass estimators are used for low tracer numbers, but these have been shown to be potentially quite biased (2019MNRAS.484..245L). More sophisticated, unbiased, mass-modeling methods for low tracer numbers are only just being developed (e.g. 2025arXiv250303812N).
@@
Secondly, there is a well-known degeneracy between the cumulative mass (which we are typically most interested in) and the velocity anisotropy profile – a measure of how circular or radial the orbit distribution is (e.g. 1987MNRAS.224...13D; 1987ApJ...313..121M; 1990AJ.....99.1548M; Wilkinson:2001ut; 2003MNRAS.343..401L; 2009MNRAS.395...76D; 2017MNRAS.471.4541R; 2021MNRAS.501..978R). For Jeans modeling, this occurs because the Jeans equations do not constitute a closed system, meaning that for a given mass model and tracer profile, a unique solution is not guaranteed (e.g. Binney2008). This typically arises when only projected line-of-sight (LOS) velocities are available.
@@
Two key methods have been proposed to break the mass-anisotropy degeneracy. Firstly, one can incorporate higher moments of the velocity distribution as data constraints. At infinite order, these fully specify the phase-space distribution function and ensure its positivity (1990AJ.....99.1548M; 1992ApJ...391..531D; 2013MNRAS.432.3361R). After the usually-used second moment, the next non-trivial moment for a spherical, non-rotating, pressure-supported system is the fourth moment. This encodes information about deviations from Gaussianity, including the ‘tailedness’ of the distribution and the kurtosis. It can be incorporated directly into the Jeans equations, thereby introducing a new fourth-order anisotropy parameter (1990AJ.....99.1548M; 2003MNRAS.343..401L; 2005MNRAS.363..918L; 2013MNRAS.432.3361R; 2024arXiv240412671W). Alternatively, one can perform luminosity-weighted integrals over the fourth moment to derive ‘Virial Shape Parameters’ (VSPs) that marginalize out the higher-order anisotropy (1990AJ.....99.1548M; 2014MNRAS.441.1584R; 2017MNRAS.471.4541R). Secondly, one can include proper motions (PMs) to gain additional phase-space information that is independent of the line-of-sight velocities (e.g. van2010new; 2013MNRAS.436.2598W). Until recently, such PMs have only been available for GCs (e.g. 2003ApJ...595..187M; anderson2010new; 2022MNRAS.514..806V; 2023MNRAS.522.5320D; 2024Natur.631..285H; 2025A&A...693A.104B). However, data are becoming increasingly available also for nearby dwarf galaxies, with recent mass models incorporating PMs having been recently reported for the Draco and Sculptor dSphs (2024ApJ...970....1V; 2025arXiv250820711V).
@@
Although Jeans mass-modeling methods typically use data binning (e.g. 2017MNRAS.471.4541R), this is not actually a requirement. Several authors have explored assuming a local Gaussian distribution function constrained by the second moment to create a bin-free Jeans method (e.g. 2013MNRAS.436.2598W; 2013MNRAS.429.3079M). With the widespread availability of individually resolved stellar velocities, this approach is preferable and has all the advantages of having $f$ that distribution function methods enjoy. [^2] This is true, up to a given order of moments, provided that the Jeans method has those moments self-consistently implemented in the velocity distribution function.
@@
In this work, we present the GravSphere2 mass-modeling method. This builds on GravSphere(2017MNRAS.471.4541R; 2021MNRAS.505.5686C) with the specific goals of: (i) reducing bias; (ii) maximizing the information content in the data; (iii) avoiding the need for data binning; and (iv) being able to successfully model, without bias, very low numbers of tracers. To achieve these goals, we self-consistently solve the second and fourth-order Jeans equations, following the framework in 2013MNRAS.432.3361R and extending it to include a new observable: fourth-order PMs. We show that this breaks degeneracies in the higher-order Jeans equations by solving along their three projected components. Unlike previous approaches, which have made more restrictive assumptions (e.g. 1990AJ.....99.1548M; 2003MNRAS.343..401L; 2005MNRAS.363..918L; 2013MNRAS.432.3361R; 2024arXiv240412671W), we also include fully independent and radially varying anisotropies at both orders. We model individual stellar velocities, rather than binning the data, by introducing a local non-Gaussian velocity probability distribution function (PDF) constrained by the second and fourth moments. We also model the positions of stars without binning, where data are available. We test these improvements using mock data for dSphs and we make the Python package, GravSphere2, publicly available at: https://github.com/dadams42/GravSphere2.
@@
This paper is organized as follows. In Sec. 2, we introduce the modeling methodology behind GravSphere2, including our method for solving the fourth-order Jeans equations, which we apply following 2013MNRAS.432.3361R and extend to include PMs (Sec. 2.1), its implementation using non-Gaussian local velocity PDFs for individual tracers, following 2020MNRAS.499.5806S (Sec. 2.2), and the functional forms we assume for our mass model and the second and fourth-order stellar velocity anisotropy profiles (Sec. 2.3). In Secs. 3.1 and 3.2, we introduce the mock data which we use to test GravSphere2. In Sec. 3.3, we describe how we integrate the approach introduced in Sec. 2.1 to fit these data, specifying the form of the likelihood function and our assumed priors. Sec. 4 shows the results of our tests of GravSphere2 on the mock data. We compare these results with previous methods that have been applied to these same data in the literature, including previous versions of GravSphere. We also explore the limit of low tracer numbers (i.e. corresponding to the faintest or most distant systems) and we compare the performance of GravSphere2 with simple mass estimators that are widely used in the literature. In Sec. 5, we briefly discuss the significance of this work and we add some concluding remarks, highlighting some upcoming works that use the new GravSphere2 method.

### v3

# GravSphere2: A higher order Jeans method for mass modeling spherical stellar systems

@@ Abstract
Aims. Mass-modeling methods are used to infer the gravitational field of stellar systems, from globular clusters to giant elliptical galaxies. While many methods already exist, most require assumptions on the form of the underlying distribution function or binning the data, leading to some loss of information. Furthermore, when only line-of-sight (LOS) data are available, many methods suffer from the well-known mass-anisotropy degeneracy. To overcome these limitations, we developed a new and publicly available mass modeling method, GravSphere2. It combines individual stellar velocities from LOS and proper motion (PM) measurements to solve the Jeans equations up to fourth order, without any data binning. Using flexible functional forms for the velocity anisotropy profiles at second and fourth order, we show how including additional constraints from a new observable, fourth-order PMs, allows us to obtain a full solution along the three dimensions and breaking the mass-anisotropy degeneracy at all orders. We tested our method on mock data for dwarf galaxies, showing how GravSphere2 improves on previous methods.
Methods. GravSphere2 introduces four key improvements over previous Jeans mass modeling methods in the literature: (i) we included fourth-order velocity moment equations in both the LOS and PM directions, for the first time, using them to break model degeneracies; (ii) we used a fully general treatment of both the second and fourth-order velocity anisotropies; (iii) we introduced a “bin-free” approach where we fit individual tracer velocities and positions using flexible and self-consistent probability density functions that include kurtosis; and (iv) we improved the likelihood sampling by using the nested sampler dynesty.
Results. GravSphere2 was able to recover the mass density, stellar velocity anisotropy, and the logarithmic slope of the mass density profile within its quoted 95% confidence intervals across almost all mocks over a wide radial range ($0.1\lesssim r/R_{1/2}\lesssim 10$, where $R_{1/2}$ is the projected half-light radius). As the number of tracers is lowered (even down to just ten tracers) it gracefully degrades, with larger uncertainties but no induced bias. We find that GravSphere2 outperforms simple mass estimators, suggesting that it is worth using even when only a few LOS velocities are available. Using 1,000 tracers without PMs, GravSphere2 recovers the logarithmic density slope at $R_{1/2}$ with $12\%$ (25%) statistical errors for cuspy (cored) mock data, enabling us to make a distinction between the two. When including PMs, this result can be improved to $8\%$ (12%). With only 100 tracers and no PMs, we were still able to recover slopes with $\sim 30\%$ (20%) errors. GravSphere2 will become a valuable new tool to hunt for massive black holes and invisible dark matter in spherical stellar systems, from globular clusters and dwarf galaxies to giant ellipticals and galaxy clusters.
@@ 1 Introduction
Many stellar systems are approximately spherical, including globular clusters (GCs) and nearby dwarf spheroidal galaxies (dSphs). Both GCs and dSphs contain seemingly invisible mass components that can be inferred from the anomalous motions of their stars and gas (see e.g., Bosma 1978; Rubin et al. 1980; Trimble 1987; Łokas 2002; Van der Marel and Anderson 2010; Battaglia et al. 2013, for more recent examples in GCs and dSphs; see e.g., Evans et al. 2022; Dickson et al. 2024; Smith et al. 2024a; Bañares-Hernández et al. 2025 and Read et al. 2019; Hayashi et al. 2020; Battaglia and Nipoti 2022; Bañares-Hernández et al. 2023; De Leo et al. 2024; Arroyo-Polonio et al. 2025; Pascale et al. 2025; Vitral et al. 2025). For GCs, the invisible mass most likely constitutes stellar remnants or an intermediate-mass black hole (IMBH), although some GCs may also contain dark matter. By contrast, dSphs appear to be dominated at all radii by dark matter. With increasingly available spectroscopic and astrometric data from recent surveys, GCs promise an unprecedented natural laboratory to study stellar remnants and hunt for IMBHs (Mezcua, 2017; Neumayer et al., 2020; Greene et al., 2020; Askar et al., 2023; Vitral et al., 2023; Häberle et al., 2024; Bañares-Hernández et al., 2025), while dSphs promise the same for dark matter (Boehm et al., 2017; Read et al., 2018, 2019; Battaglia and Nipoti, 2022; De Leo et al., 2024; Vitral et al., 2024, 2025; Zimmermann et al., 2025).
The above proliferation in data has led to an associated proliferation in mass modeling methodologies. All methods seek to solve the steady-state Collisionless Boltzmann Equation (CBE), also known as the Vlasov Equation (e.g., Binney and Tremaine 2008), expressed as
@@
where $\Phi$ is the gravitational potential resulting from all mass components in the system. Then, satisfying the Poisson equation with mass density, $\rho$, we have
@@
By solving for the gravitational potential and subtracting the expected potential from all visible components we can infer the presence of invisible masses for future study, such as black holes and dark matter (e.g., Van der Marel and Anderson 2010; Walker et al. 2009), which is usually done in the steady state approximation, assuming $\frac{\partial f}{\partial t}=0$.
@@
There are number of methods for solving the CBE. Jeans methods solve moments of the CBE (Jeans, 1922), having the advantages that they are fast [^1] This is particularly timely when assessing increasingly large samples with thousands (in some cases approaching millions, Häberle et al. 2025) of tracers. and minimize potential biases that could creep in due to assumptions about the form of the distribution function, $f$ (e.g., Cappellari 2008; Watkins et al. 2013; Mamon et al. 2013; Cappellari 2015; Read and Steger 2017). The main disadvantages, however, are that the data typically need to be binned, while $f$ is not determined. This leads to some loss of information and makes it hard to incorporate stars of uncertain membership, binaries, and/or rigorously model selection effects (e.g. Read et al., 2021). Furthermore, since $f$ is not determined, a physical $f$ is not guaranteed. This can result in models that formally require a negative $f$ in some regions of phase space (Dejonghe and Merritt, 1992; Richardson and Fairbairn, 2013) which should then be discarded (e.g., An and Evans, 2006).
@@
Distribution function methods assume some form for $f$ and then fit this to the data (e.g., Gieles and Zocchi 2015; Vasiliev 2019; Read et al. 2021; Dickson et al. 2023). In such cases, $f$ can also be implicitly expressed as a superposition of orbits with Schwarzschild methods (e.g., Schwarzschild 1979; Chanamé et al. 2008; van den Bosch et al. 2008; Breddels et al. 2013) or with made-to-measure methods, in which $N$-body simulations are evolved to fit some data constraints (e.g., Syer and Tremaine 1996; de Lorenzi et al. 2007; Dehnen 2009; Hunt and Kawata 2013). In addition, Li et al. (2025) proposed a novel nonparametric distribution function method. Distribution function methods have the advantages that no data binning is required and $f$ is determined by the data. However, if the functional form for $f$ (or its implicit specification via other methods) does not encompass the true solution, then bias can creep in.
@@
Finally, numerous machine-learning methods have been proposed that learn from training data to either accelerate distribution function fitting or incorporate cosmological priors into the mass modeling pipeline (e.g., Expósito-Márquez et al. 2023; Nguyen et al. 2025; Lim et al. 2025; Sarrato-Alós et al. 2025).
@@
Given the proliferation of mass modeling methods, we may wonder why yet another method is needed. However, there are two key problems that remain to be addressed, even for spherical stellar systems. Firstly, while the amount and quality of data continue to improve, there will always be interesting systems with very few stellar tracers (e.g., Collins et al. 2021; Smith et al. 2024b; Pickett et al. 2025). Typically, simple mass estimators are used for low tracer numbers, but these have been shown to be potentially significantly biased (Laporte et al., 2019). More sophisticated, unbiased, mass modeling methods for low tracer numbers are only just being developed (e.g., Nguyen et al. 2025).
@@
Secondly, there is a well-known degeneracy between the cumulative mass (which we are typically most interested in) and the velocity anisotropy profile, which is a measure of how circular or radial the orbit distribution is (e.g., Dejonghe 1987; Merritt 1987; Merrifield and Kent 1990; Wilkinson et al. 2002; Łokas and Mamon 2003; de Lorenzi et al. 2009; Read and Steger 2017; Read et al. 2021). For Jeans modeling, this occurs because the Jeans equations do not constitute a closed system, meaning that for a given mass model and tracer profile, a unique solution is not guaranteed (e.g., Binney and Tremaine 2008). This typically arises when only projected line-of-sight (LOS) velocities are available.
@@
Two key methods have been proposed to break the mass-anisotropy degeneracy. Firstly, we can incorporate higher moments of the velocity distribution as data constraints. At infinite order, these fully specify the phase-space distribution function and ensure its positivity (Merrifield and Kent, 1990; Dejonghe and Merritt, 1992; Richardson and Fairbairn, 2013). After the usually-used second moment, the next nontrivial moment for a spherical, nonrotating, pressure-supported system is the fourth moment. This encodes information about deviations from Gaussianity, including the ”tailedness” of the distribution and the kurtosis. It can be incorporated directly into the Jeans equations, thereby introducing a new fourth-order anisotropy parameter (Merrifield and Kent, 1990; Łokas and Mamon, 2003; Łokas et al., 2005; Richardson and Fairbairn, 2013; Wardana et al., 2025). Alternatively, we can perform luminosity-weighted integrals over the fourth moment to derive the virial shape parameters (VSPs) that marginalize out the higher-order anisotropy (Merrifield and Kent, 1990; Richardson and Fairbairn, 2014; Read and Steger, 2017). Secondly, we can include proper motions (PMs) to gain additional phase-space information that is independent of the LOS velocities (e.g., Van der Marel and Anderson 2010; Watkins et al. 2013). Until recently, such PMs have only been available for GCs (e.g., McNamara et al. 2003; Anderson and Van der Marel 2010; Vitral et al. 2022; Dickson et al. 2023; Häberle et al. 2024; Bañares-Hernández et al. 2025). However, data are becoming increasingly available for nearby dwarf galaxies, with recent mass models incorporating PMs recently reported for the Draco and Sculptor dSphs (Vitral et al., 2024, 2025).
@@
Although Jeans mass modeling methods typically use data binning (e.g. Read and Steger, 2017), this is not actually a requirement. Several authors have explored assuming a local Gaussian distribution function constrained by the second moment to create a bin-free Jeans method (e.g., Watkins et al. 2013; Mamon et al. 2013). With the widespread availability of individually resolved stellar velocities, this approach is preferable and has all the advantages of having $f$ that distribution function methods enjoy. [^2] This is true, up to a given order of moments, provided that the Jeans method has those moments self-consistently implemented in the velocity distribution function.
@@
In this work, we present the GravSphere2 mass modeling method. This builds on GravSphere(Read and Steger, 2017; Collins et al., 2021) with the specific goals of: (i) reducing bias; (ii) maximizing the information content in the data; (iii) avoiding the need for data binning; and (iv) having the capacity able to successfully model, without bias, very low numbers of tracers. To achieve these goals, we self-consistently solved the second- and fourth-order Jeans equations, following the framework in Richardson and Fairbairn (2013), extending it to include a new observable: fourth-order PMs. We show that this approach breaks the degeneracies in the higher order Jeans equations by solving along their three projected components. Unlike previous approaches, which have made more restrictive assumptions (e.g., Merrifield and Kent, 1990; Łokas and Mamon, 2003; Łokas et al., 2005; Richardson and Fairbairn, 2013; Wardana et al., 2025), we also included fully independent and radially varying anisotropies at both orders. We modeled the individual stellar velocities, rather than binning the data, by introducing a local non-Gaussian velocity probability distribution function (PDF) constrained by the second and fourth moments. We also modeled the positions of stars without binning, when the relevant data were available. We tested these improvements using mock data for dSphs and put together the Python package, GravSphere2, which is now publicly available [^3] Available at: https://github.com/dadams42/GravSphere2.
@@
This paper is organized as follows. In Sect. 2, we introduce the modeling methodology behind GravSphere2, including our method for solving the fourth-order Jeans equations, which we applied following Richardson and Fairbairn (2013) and extended to include PMs (Sect. 2.1). We also describe its implementation using non-Gaussian local velocity PDFs for individual tracers, following Sanders and Evans 2020 (Sect. 2.2), and the functional forms we assume for our mass model and the second and fourth-order stellar velocity anisotropy profiles (Sect. 2.3). In Sects. 3.1 and 3.2, we introduce the mock data we used to test GravSphere2. In Sect. 3.3, we describe how we integrate the approach introduced in Sect. 2.1 to fit these data, specifying the form of the likelihood function and our assumed priors. Sect. 4 shows the results of our tests of GravSphere2 on the mock data. We compared these results with previous methods that were applied to these same data in the literature, including previous versions of GravSphere. We also explore the limit of low tracer numbers (i.e., corresponding to the faintest or most distant systems) and we compare the performance of GravSphere2 with simple mass estimators that are widely used in the literature. In Sect. 5, we briefly discuss the significance of this work and add our concluding remarks, highlighting upcoming works that make use of the new GravSphere2 method.